OpenStudy (anonymous):

Consider the following definition. Definition. An integer n is sane if 3 | (n2 + 2n). Give a counterexample to the following: All odd integers are sane. n =

OpenStudy (unklerhaukus):

what do you get for n =1, n=3, n=5 ?

OpenStudy (unklerhaukus):

for n=1 n^2 +2n --> (1)^2+2(1) = 1+2 = 3 which is a multiple of 3, so n=1 in not a counter example

OpenStudy (anonymous):

incorrect

OpenStudy (unklerhaukus):

for n=3 n^2 + 2n --> 3^2 + 2(3) =

OpenStudy (anonymous):

same, I wonder if I'm going to have to have the problem written out like the original one is

OpenStudy (unklerhaukus):

what do you get for the n=3 case?

OpenStudy (unklerhaukus):

@Scaleman

OpenStudy (anonymous):

incorrect

OpenStudy (unklerhaukus):

what about the next smallest odd integer: n=5 ?

OpenStudy (unklerhaukus):

5^2+2(5) = ?

OpenStudy (unklerhaukus):

@Scaleman ?

OpenStudy (unklerhaukus):

is 5 sane?

OpenStudy (anonymous):

Thank you very much for the help